OpenStudy (anonymous):

mr. smith's tax on his house for 1965 was $283.79. if the tax rate for that year was $3.835 per $100 of assessed valuation, for what amount was mr. smith's house assessed? A. $10.88 B. $74 C. $740 D. $1088.33 E. $7400

OpenStudy (anonymous):

Use unitary method. tax is $3.85, for every $100 of value for $1 tax, value is $100/ 3.85 So for $283.79 tax the value would be $283.79* 100/3.85 = ____________Solve for the Answer

OpenStudy (anonymous):

did you understand the method?

OpenStudy (mathstudent55):

Use a proportion. If for $100 of value, tax is $3.835, then for x value, tax is $283.79. In proportion "talk" this can be stated as: $100 is to $3.835 as x is to $283.79 \( \dfrac{$100}{$3.835} = \dfrac{x}{$283.79} \) Cross multiply and solve for x.